# multiMentalRoBERTa: A Fine-tuned Multiclass Classifier for Mental Health Disorder

### v1

@@ I Introduction
The World Health Organization defines mental health as “a state of mental well-being that enables people to cope with the stresses of life, realize their abilities, learn and work well, and contribute to their community,” and of mental disorders as clinically significant disturbances in cognition, emotional regulation, or behavior that cause distress or functional impairment[1]. World Health Organization Mental disorders are common: In 2019, approximately one in eight people worldwide (970 million) lived with a mental disorder, with anxiety and depressive disorders the most prevalent; early pandemic estimates showed sharp increases in both conditions [2]. The public-health burden is underscored by suicide, which claims more than 720,000 lives annually and ranks as a leading cause of death among young people (ages 15 - 29) globally[3].
@@
In modern times, many individuals turn to social media platforms and peer-support forums to share experiences and seek help for mental health concerns. Among the most commonly discussed conditions are depression (characterized by persistent sadness, hopelessness, and loss of interest in daily activities [2]), anxiety disorders (involving excessive fear or worry that interferes with daily functioning [2]) and post-traumatic stress disorder (PTSD) (marked by intrusive memories, avoidance, and hyperarousal following traumatic events [4]). Another frequent concern is suicidal ideation, defined as persistent thoughts about or planning of taking one’s own life [3]. Stress is also widely examined in online contexts, though it is not formally classified as a mental disorder. Instead, it is a psychological state that can contribute to or worsen mental illnesses, with the National Center for Complementary and Integrative Health (NCCIH) noting that long-term stress may increase vulnerability to conditions such as depression and anxiety [5]. In our dataset, we also observe that stress occupies an intermediate space between mental health disorders and neutral data (see Figure 1).
@@
Classifying mental health disorders in peer-support systems and online forums is critical, as these platforms are often where individuals first disclose symptoms due to accessibility and anonymity. Accurate classification enables timely support, risk detection, and referral to appropriate resources, which is particularly important for conditions such as depression and suicidal ideation. Advances in artificial intelligence (AI) and machine learning (ML) provide effective tools for this task, as they can analyze large volumes of social media data, identify linguistic and behavioral patterns, and improve early detection of mental health problems [6]. Beyond scalability, AI has shown promise in enhancing diagnostic accuracy and addressing global shortages of mental health professionals, offering a cost-effective way to expand access to care [7]. A key difficulty arises from the contrast between binary and multi-class classification: in binary setups, random guessing yields a 50% accuracy, whereas in a six-class problem, it drops to about 16.7% (1/6). This gap underscores the heightened complexity of multi-class classification, where models face a more demanding decision space with a narrower margin for error [8]. Building on this context, the present work contributes to advancing multiclass detection of mental health disorders through the following aims:
@@ II-A Peer Support and Social Media in Mental Health Research
Online peer-support platforms and social media communities have become important spaces for individuals to share experiences, seek advice, and disclose symptoms related to mental health. Platforms such as Reddit, Twitter, and Facebook not only facilitate supportive exchanges but also provide rich data sources for analyzing depression, stress, anxiety, PTSD, and suicidality [9]. These communities help reduce stigma, encourage help-seeking, and often provide early signals of distress, complementing traditional clinical pathways. Self-disclosure plays a central role, as individuals candidly share vulnerable experiences. Research shows that self-disclosure supports psychological well-being and therapeutic progress [10], and studies of Reddit forums indicate that highly self-disclosing posts, though often receiving fewer upvotes, attract faster and more extensive responses, reflecting strong community support for vulnerable users.
@@ II-B Computational Approaches for Mental Health Disorder Detection
Early computational approaches to mental health detection from non-clinical texts relied on traditional machine learning methods such as Support Vector Machines (SVM), logistic regression, and decision trees, using lexical features (e.g., bag-of-words, TF-IDF), psycholinguistic lexicons like LIWC, and basic behavioral indicators. These methods showed feasibility but struggled with generalization and capturing nuanced semantics [11]. With deep learning, research shifted to CNNs, RNNs, and attention-based models. Ji et al. [12] introduced an Attentive Relation Network combining sentiment lexicons, latent topics, and embeddings to model relationships between suicidal ideation and related disorders, outperforming strong baselines across datasets. Subsequent work advanced the field with deep contextualized representations. Jiang et al. [13] introduced a large Reddit-based dataset spanning eight psychiatric disorders and showed that BERT-based models substantially outperformed LIWC baselines, with user-level modeling proving more effective than post-level classification. Still, multiclass classification remained difficult due to overlap and comorbidity. More recent studies explored multimodal and knowledge-augmented models: Toto et al. [14] proposed AudiBERT, integrating audio and text with dual self-attention for depression screening, while Yang et al. [15] developed KC-Net, which combines commonsense mental state knowledge with supervised contrastive learning for stress and depression detection.
@@ II-C Explainability in Mental Health Models
The application of AI to mental health detection requires not only high predictive accuracy but also explainability, as clinicians, peer support specialists, and users must trust model outputs in high-stakes contexts. Recent surveys distinguish between explainable AI (XAI), which provides post-hoc explanations, and interpretable AI (IAI), where models are inherently transparent [16]. Most healthcare NLP work emphasizes local post-hoc methods such as LIME, SHAP, and attention visualization, while global interpretability remains limited. In mental health, explainability is increasingly linked to clinical grounding: Bao et al. [17] showed that transformer-based models can detect depressive symptoms and generate explanations aligned with clinical assessments (e.g., PHQ-9), with extractive explanations proving more reliable than abstractive ones. Joyce et al. [18] similarly stressed the need for understandability - combining transparency and interpretability - through their TIFU framework, which aligns explanations with psychiatric reasoning.
@@ II-D Advancing Mental Health Classification
Recent work such as MentaLLaMA[19] has advanced the field by introducing an instruction-tuned large language model for interpretable mental health analysis. Its major focus is on text generation, producing both classifications and explanatory rationales in a single step. To achieve this, MentaLLaMA requires fine-tuning LLaMA2 models with 7B or 13B parameters on the IMHI dataset. While it demonstrates strong generalization and generates human-readable explanations, its evaluation also showed that discriminative models like MentalBERT[20] still achieve superior classification accuracy across several benchmarks. MentalBERT addressed the gap of lacking domain-specific pretrained language models by continuing pretraining on large-scale Reddit corpora from mental health communities. The resulting models improved performance over general-purpose BERT and RoBERTa and provided the first publicly available pretrained embeddings for the domain. To be effectively applied, however, these embeddings still require fine-tuning on specific downstream classification tasks. Most evaluations focused on binary detection tasks such as depression vs. control or suicidal vs. non-suicidal, with only limited exploration of multiclass classification. In addition, the work did not explicitly address explainability or the correlation of overlapping disorders.
@@ III-A Data Collection & Curation
Reddit Depression Dataset[21] - is constructed from multiple balanced subsets of Reddit and forum posts, with 400 documents per class. Labels are not human-annotated but are assigned automatically based on subreddit membership. Posts from the Depression Support subreddit (DS) are treated as depression cases, while additional subsets such as posts from users with self-declared depression in other subreddits (DO) are also included. Control subsets from subreddits like Breast Cancer (BC), Family/Friendship (FF), and non-depression forums (DND) are provided in the original dataset but are not used in this research. For the present study, only the DS and DO subsets are selected, as they provide self-disclosure and broader depression-related content most relevant to the research objective.
Dreaddit Dataset[22] - is constructed from Reddit posts collected between January 2017 and November 2018 across multiple subreddits. It is human-annotated through Amazon Mechanical Turk, where each sampled segment is labeled for the presence of stress. For this research, only posts related to stress, anxiety, and PTSD are collected. Stress data are obtained from subreddits focused on abuse, financial need, and social relationships, while anxiety posts are taken from r/anxiety, and PTSD posts are taken from r/ptsd. The dataset therefore provides human-labeled user-generated text that captures mental health challenges across these three categories, which are retained for analysis.
The Stress Annotated Dataset (SAD) [23] - is composed of 6,850 SMS-like sentences that are generated through chatbot interactions, crowdsourcing, and targeted web scraping from an online repository. Each sentence is labeled by multiple annotators with one or two stressor categories and a severity rating, and nine final categories are established, including work, school, financial problems, family issues, health or fatigue, social relationships, emotional turmoil, everyday decision making, and other. For this research, only the stress-labeled sentences are retained to ensure that the dataset is focused exclusively on texts that express stressful experiences.
Reddit Suicide and Depression (SDCNL) Dataset [24] - is constructed by collecting 1,895 posts from two mental health subreddits, where posts from r/SuicideWatch are labeled as suicidal and posts from r/Depression are labeled as depressed. Each post is retained with its original text and subreddit-based label, although the annotations are inherently noisy due to reliance on self-reporting. For this research, only the suicidal and depression data are selected to enable a focused analysis on distinguishing between depressive expression and suicidal ideation.
@@
The use of social media data in mental health research also raises ethical challenges. While posts are technically public, users often perceive them as semi-private, with limited awareness of data permanence and computational analysis [25]. Ethical guidelines emphasize informed consent where possible, de-identification, secure storage, and sensitive handling of disclosures. Scholars recommend practices such as paraphrasing examples and avoiding cross-platform linkages that could re-identify individuals. Importantly, for this work no user-identifiable information is utilized, and the classification models developed are not capable of exposing sensitive information, thereby aligning with these ethical principles.
@@ III-B Data Exploration
Figure 1 provides a visual representation of how different mental health conditions are positioned in the embedding space, and the patterns observed are consistent with the clustering and correlation results. Table III shows how true labels are distributed across clusters when RoBERTa embeddings are applied, while Table IV quantifies correlations among the classes. The visualization and tables together indicate that None is concentrated almost entirely in a single cluster, suggesting strong separability, whereas Stress and Depression are spread across multiple clusters. Notably, Suicidal and PTSD overlap with Depression, reflecting linguistic similarity. The correlation matrix confirms that the highest correlation (0.929) occurs between Depression and Suicidal, followed by strong associations between Anxiety and PTSD (0.843) and between Stress and PTSD (0.748). In contrast, the None class shows very low correlation with all other categories, reinforcing its distinct nature. The widespread distribution of Stress across clusters can be explained by its role as a broad underlying factor that contributes to multiple psychiatric conditions. Prior work [26] highlights that chronic stress is strongly associated with the onset of major depressive disorder, PTSD, and other serious mental illnesses, which supports the observation that stress does not remain isolated but overlaps with other mental health categories. These overlaps may be attributed to shared vocabulary, co-occurring symptoms, and thematic similarities across conditions, which can blur class boundaries.
@@ III-D Transformer Fine-Tuned Models
MentalBERT [^5] https://huggingface.co/mental/mental-bert-base-uncased (mental-roberta-base [^6] https://huggingface.co/mental/mental-roberta-base) - is a domain-adapted BERT model pretrained specifically on large-scale mental health-related Reddit data. Unlike general-purpose BERT and RoBERTa, mentalBERT’s pretraining corpus is drawn from online communities where users disclose experiences of depression, stress, and suicidal ideation. This makes its language representations closely aligned with the discourse and symptom expressions typical of peer-support forums. Prior work has shown that such domain-specific pretraining improves robustness and reliability, as the model learns subtle linguistic markers, colloquial expressions, and context-specific cues that general models often overlook [27]. In this study, we fine-tune the best-performing variant, mental/mental-roberta-base, to evaluate whether leveraging both large-scale pretraining and mental health–oriented domain adaptation can further enhance classification performance across depression, anxiety, PTSD, stress, suicidal ideation, and neutral content.
@@ III-E Prompt Engineering
Large language models are evaluated through structured prompts designed to guide classification of mental health texts. Each prompt begins with a clear task definition, contextual notes to prevent medical advice or crisis instructions, and an explicit list of predefined labels. To reduce ambiguity, rules for handling uncertainty are also included, instructing the model to choose the closest category or return none when no signal of mental health content is present. In addition, the models are explicitly instructed to assign an Unknown label when the input cannot be confidently classified, a strategy shown to reduce hallucinations and improve reliability in prior work [28]. The output format is constrained to a strict schema (id - predicted_label - text) to ensure reproducible and interpretable evaluation.
@@
All prompting experiments are executed through AI agents (GPT [^8] https://openai.com/api/ and DeepSeek [^9] https://api-docs.deepseek.com/) using API endpoints. Generation parameters are fixed with a batch size of 5, temperature of 0.0, and top-p of 1.0 to encourage deterministic outputs while permitting limited diversity [29]. An exception is GPT-5, which currently supports only a fixed temperature of 1.0 (see Code Listing 1), resulting in less constrained outputs. Together, these prompting strategies and parameter configurations enable systematic evaluation of model reasoning without task-specific fine-tuning.

### v2

@@ I Introduction
The World Health Organization defines mental health as “a state of mental well-being that enables people to cope with the stresses of life, realize their abilities, learn and work well, and contribute to their community,” and of mental disorders as clinically significant disturbances in cognition, emotional regulation, or behavior that cause distress or functional impairment[Mental-health]. World Health Organization Mental disorders are common: In 2019, approximately one in eight people worldwide (970 million) lived with a mental disorder, with anxiety and depressive disorders the most prevalent; early pandemic estimates showed sharp increases in both conditions [Mental-disorders]. The public-health burden is underscored by suicide, which claims more than 720,000 lives annually and ranks as a leading cause of death among young people (ages 15 - 29) globally[Suicide].
@@
In modern times, many individuals turn to social media platforms and peer-support forums to share experiences and seek help for mental health concerns. Among the most commonly discussed conditions are depression (characterized by persistent sadness, hopelessness, and loss of interest in daily activities [Mental-disorders]), anxiety disorders (involving excessive fear or worry that interferes with daily functioning [Mental-disorders]) and post-traumatic stress disorder (PTSD) (marked by intrusive memories, avoidance, and hyperarousal following traumatic events [PTSD]). Another frequent concern is suicidal ideation, defined as persistent thoughts about or planning of taking one’s own life [Suicide]. Stress is also widely examined in online contexts, though it is not formally classified as a mental disorder. Instead, it is a psychological state that can contribute to or worsen mental illnesses, with the National Center for Complementary and Integrative Health (NCCIH) noting that long-term stress may increase vulnerability to conditions such as depression and anxiety [Stress]. In our dataset, we also observe that stress occupies an intermediate space between mental health disorders and neutral data (see Figure 1).
@@
Classifying mental health disorders in peer-support systems and online forums is critical, as these platforms are often where individuals first disclose symptoms due to accessibility and anonymity. Accurate classification enables timely support, risk detection, and referral to appropriate resources, which is particularly important for conditions such as depression and suicidal ideation. Advances in artificial intelligence (AI) and machine learning (ML) provide effective tools for this task, as they can analyze large volumes of social media data, identify linguistic and behavioral patterns, and improve early detection of mental health problems [abd2020application]. Beyond scalability, AI has shown promise in enhancing diagnostic accuracy and addressing global shortages of mental health professionals, offering a cost-effective way to expand access to care [minerva2023ai]. A key difficulty arises from the contrast between binary and multi-class classification: in binary setups, random guessing yields a 50% accuracy, whereas in a six-class problem, it drops to about 16.7% (1/6). This gap underscores the heightened complexity of multi-class classification, where models face a more demanding decision space with a narrower margin for error [fields2024integrating]. Building on this context, the present work contributes to advancing multiclass detection of mental health disorders through the following aims:
@@ II-A Peer Support and Social Media in Mental Health Research
Online peer-support platforms and social media communities have become important spaces for individuals to share experiences, seek advice, and disclose symptoms related to mental health. Platforms such as Reddit, Twitter, and Facebook not only facilitate supportive exchanges but also provide rich data sources for analyzing depression, stress, anxiety, PTSD, and suicidality [garg2023mental]. These communities help reduce stigma, encourage help-seeking, and often provide early signals of distress, complementing traditional clinical pathways. Self-disclosure plays a central role, as individuals candidly share vulnerable experiences. Research shows that self-disclosure supports psychological well-being and therapeutic progress [balani2015detecting], and studies of Reddit forums indicate that highly self-disclosing posts, though often receiving fewer upvotes, attract faster and more extensive responses, reflecting strong community support for vulnerable users.
@@ II-B Computational Approaches for Mental Health Disorder Detection
Early computational approaches to mental health detection from non-clinical texts relied on traditional machine learning methods such as Support Vector Machines (SVM), logistic regression, and decision trees, using lexical features (e.g., bag-of-words, TF-IDF), psycholinguistic lexicons like LIWC, and basic behavioral indicators. These methods showed feasibility but struggled with generalization and capturing nuanced semantics [calvo2017natural]. With deep learning, research shifted to CNNs, RNNs, and attention-based models. Ji et al. [ji2022suicidal] introduced an Attentive Relation Network combining sentiment lexicons, latent topics, and embeddings to model relationships between suicidal ideation and related disorders, outperforming strong baselines across datasets. Subsequent work advanced the field with deep contextualized representations. Jiang et al. [jiang2020detection] introduced a large Reddit-based dataset spanning eight psychiatric disorders and showed that BERT-based models substantially outperformed LIWC baselines, with user-level modeling proving more effective than post-level classification. Still, multiclass classification remained difficult due to overlap and comorbidity. More recent studies explored multimodal and knowledge-augmented models: Toto et al. [toto2021audibert] proposed AudiBERT, integrating audio and text with dual self-attention for depression screening, while Yang et al. [yang2022mental] developed KC-Net, which combines commonsense mental state knowledge with supervised contrastive learning for stress and depression detection.
@@ II-C Explainability in Mental Health Models
The application of AI to mental health detection requires not only high predictive accuracy but also explainability, as clinicians, peer support specialists, and users must trust model outputs in high-stakes contexts. Recent surveys distinguish between explainable AI (XAI), which provides post-hoc explanations, and interpretable AI (IAI), where models are inherently transparent [huang2024explainable]. Most healthcare NLP work emphasizes local post-hoc methods such as LIME, SHAP, and attention visualization, while global interpretability remains limited. In mental health, explainability is increasingly linked to clinical grounding: Bao et al. [bao2024explainable] showed that transformer-based models can detect depressive symptoms and generate explanations aligned with clinical assessments (e.g., PHQ-9), with extractive explanations proving more reliable than abstractive ones. Joyce et al. [joyce2023explainable] similarly stressed the need for understandability - combining transparency and interpretability - through their TIFU framework, which aligns explanations with psychiatric reasoning.
@@ II-D Advancing Mental Health Classification
Recent work such as MentaLLaMA[yang2023mentalllama] has advanced the field by introducing an instruction-tuned large language model for interpretable mental health analysis. Its major focus is on text generation, producing both classifications and explanatory rationales in a single step. To achieve this, MentaLLaMA requires fine-tuning LLaMA2 models with 7B or 13B parameters on the IMHI dataset. While it demonstrates strong generalization and generates human-readable explanations, its evaluation also showed that discriminative models like MentalBERT[ji2022mentalbert] still achieve superior classification accuracy across several benchmarks. MentalBERT addressed the gap of lacking domain-specific pretrained language models by continuing pretraining on large-scale Reddit corpora from mental health communities. The resulting models improved performance over general-purpose BERT and RoBERTa and provided the first publicly available pretrained embeddings for the domain. To be effectively applied, however, these embeddings still require fine-tuning on specific downstream classification tasks. Most evaluations focused on binary detection tasks such as depression vs. control or suicidal vs. non-suicidal, with only limited exploration of multiclass classification. In addition, the work did not explicitly address explainability or the correlation of overlapping disorders.
@@ III-A Data Collection & Curation
Reddit Depression Dataset[pirina2018identifying] - is constructed from multiple balanced subsets of Reddit and forum posts, with 400 documents per class. Labels are not human-annotated but are assigned automatically based on subreddit membership. Posts from the Depression Support subreddit (DS) are treated as depression cases, while additional subsets such as posts from users with self-declared depression in other subreddits (DO) are also included. Control subsets from subreddits like Breast Cancer (BC), Family/Friendship (FF), and non-depression forums (DND) are provided in the original dataset but are not used in this research. For the present study, only the DS and DO subsets are selected, as they provide self-disclosure and broader depression-related content most relevant to the research objective.
Dreaddit Dataset[turcan2019dreaddit] - is constructed from Reddit posts collected between January 2017 and November 2018 across multiple subreddits. It is human-annotated through Amazon Mechanical Turk, where each sampled segment is labeled for the presence of stress. For this research, only posts related to stress, anxiety, and PTSD are collected. Stress data are obtained from subreddits focused on abuse, financial need, and social relationships, while anxiety posts are taken from r/anxiety, and PTSD posts are taken from r/ptsd. The dataset therefore provides human-labeled user-generated text that captures mental health challenges across these three categories, which are retained for analysis.
The Stress Annotated Dataset (SAD) [mauriello2021sad] - is composed of 6,850 SMS-like sentences that are generated through chatbot interactions, crowdsourcing, and targeted web scraping from an online repository. Each sentence is labeled by multiple annotators with one or two stressor categories and a severity rating, and nine final categories are established, including work, school, financial problems, family issues, health or fatigue, social relationships, emotional turmoil, everyday decision making, and other. For this research, only the stress-labeled sentences are retained to ensure that the dataset is focused exclusively on texts that express stressful experiences.
Reddit Suicide and Depression (SDCNL) Dataset [haque2021deep] - is constructed by collecting 1,895 posts from two mental health subreddits, where posts from r/SuicideWatch are labeled as suicidal and posts from r/Depression are labeled as depressed. Each post is retained with its original text and subreddit-based label, although the annotations are inherently noisy due to reliance on self-reporting. For this research, only the suicidal and depression data are selected to enable a focused analysis on distinguishing between depressive expression and suicidal ideation.
@@
The use of social media data in mental health research also raises ethical challenges. While posts are technically public, users often perceive them as semi-private, with limited awareness of data permanence and computational analysis [benton2017ethical]. Ethical guidelines emphasize informed consent where possible, de-identification, secure storage, and sensitive handling of disclosures. Scholars recommend practices such as paraphrasing examples and avoiding cross-platform linkages that could re-identify individuals. Importantly, for this work no user-identifiable information is utilized, and the classification models developed are not capable of exposing sensitive information, thereby aligning with these ethical principles.
@@ III-B Data Exploration
Figure 1 provides a visual representation of how different mental health conditions are positioned in the embedding space, and the patterns observed are consistent with the clustering and correlation results. Table III shows how true labels are distributed across clusters when RoBERTa embeddings are applied, while Table IV quantifies correlations among the classes. The visualization and tables together indicate that None is concentrated almost entirely in a single cluster, suggesting strong separability, whereas Stress and Depression are spread across multiple clusters. Notably, Suicidal and PTSD overlap with Depression, reflecting linguistic similarity. The correlation matrix confirms that the highest correlation (0.929) occurs between Depression and Suicidal, followed by strong associations between Anxiety and PTSD (0.843) and between Stress and PTSD (0.748). In contrast, the None class shows very low correlation with all other categories, reinforcing its distinct nature. The widespread distribution of Stress across clusters can be explained by its role as a broad underlying factor that contributes to multiple psychiatric conditions. Prior work [davis2017neurobiology] highlights that chronic stress is strongly associated with the onset of major depressive disorder, PTSD, and other serious mental illnesses, which supports the observation that stress does not remain isolated but overlaps with other mental health categories. These overlaps may be attributed to shared vocabulary, co-occurring symptoms, and thematic similarities across conditions, which can blur class boundaries.
@@ III-D Transformer Fine-Tuned Models
MentalBERT [^5] https://huggingface.co/mental/mental-bert-base-uncased (mental-roberta-base [^6] https://huggingface.co/mental/mental-roberta-base) - is a domain-adapted BERT model pretrained specifically on large-scale mental health-related Reddit data. Unlike general-purpose BERT and RoBERTa, mentalBERT’s pretraining corpus is drawn from online communities where users disclose experiences of depression, stress, and suicidal ideation. This makes its language representations closely aligned with the discourse and symptom expressions typical of peer-support forums. Prior work has shown that such domain-specific pretraining improves robustness and reliability, as the model learns subtle linguistic markers, colloquial expressions, and context-specific cues that general models often overlook [alsentzer2019publicly]. In this study, we fine-tune the best-performing variant, mental/mental-roberta-base, to evaluate whether leveraging both large-scale pretraining and mental health–oriented domain adaptation can further enhance classification performance across depression, anxiety, PTSD, stress, suicidal ideation, and neutral content.
@@ III-E Prompt Engineering
Large language models are evaluated through structured prompts designed to guide classification of mental health texts. Each prompt begins with a clear task definition, contextual notes to prevent medical advice or crisis instructions, and an explicit list of predefined labels. To reduce ambiguity, rules for handling uncertainty are also included, instructing the model to choose the closest category or return none when no signal of mental health content is present. In addition, the models are explicitly instructed to assign an Unknown label when the input cannot be confidently classified, a strategy shown to reduce hallucinations and improve reliability in prior work [islam2025llm]. The output format is constrained to a strict schema (id - predicted_label - text) to ensure reproducible and interpretable evaluation.
@@
All prompting experiments are executed through AI agents (GPT [^8] https://openai.com/api/ and DeepSeek [^9] https://api-docs.deepseek.com/) using API endpoints. Generation parameters are fixed with a batch size of 5, temperature of 0.0, and top-p of 1.0 to encourage deterministic outputs while permitting limited diversity [alammar2024hands]. An exception is GPT-5, which currently supports only a fixed temperature of 1.0 (see Code Listing 1), resulting in less constrained outputs. Together, these prompting strategies and parameter configurations enable systematic evaluation of model reasoning without task-specific fine-tuning.